# Neural density functional theory of liquid-gas phase coexistence

### v1

@@ IV.4 Outlook
As described in Sec. III.5, mean-field behaviour is found from OZ plots in the close vicinity of the critical point, inline with the construction of the neural functional with a finite box size. Nevertheless, we envisage that accessing ‘true’ critical behaviour is not excluded per se from the neural functional. The strategy used for the extrapolation to large box sizes via the input node $1/L$ might well prove to be helpful for this purpose. Supplying additional training data that feature the slow decay of density profiles associated with long-range correlations might also be necessary. Further, the architecture of the neural network might need to be modified to accommodate the growing correlation length which could require incorporating a larger density window as input to the neural functional.
Although we have focused in this work on the prototypical Lennard-Jones fluid, for ease of comparison with previous accurate simulation data, we see no fundamental problems in applying our methodology to more complex fluids. As the formal structure of DFT remains intact for arbitrary interaction potentials, the method could be transferred directly to the investigation of systems that feature more elaborate force fields, including three- and higher-body contributions, as used e.g. to model interactions in water [86, 87] or in colloidal gels [88]. For the case of anisotropic and molecular fluids, orientational degrees of freedom must be accounted for: the one-body density profile $\rho(\mathbf{r},\bm{\omega})$ which enters the functional mapping must be resolved with respect to both position $\mathbf{r}$ and orientation $\bm{\omega}$. Methods developed in the context of molecular DFT [89, 90] could serve as a practical guide on how to deal with the additional orientation-dependence. For mixtures, the relevant functionals depend on the density profiles $\rho_{i}(\mathbf{r})$ of the individual species $i=1,\dots,s$. A neural network representing the direct correlation functional $c_{1,i}(\mathbf{r};[\rho_{1},\dots,\rho_{s}],T)$ must therefore comprise the additional number of channels in the input and output layers. Tackling the fluid phase behaviour of multi-component systems is, of course, important in physical chemistry and chemical engineering. Recall, for example, that rich phase behaviour emerges already for the case of a very simple binary mixture [91].
@@ Acknowledgements.
We thank Francesco Turci and Nigel Wilding for helpful discussions. R. E. acknowledges support of the Leverhulme Trust, grant no. EM-2020-029/4.

### v2

@@ IV.4 Outlook
As described in Sec. III.5, mean-field behavior is found from OZ plots in the close vicinity of the critical point, inline with the construction of the neural functional with a finite box size. Nevertheless, we envisage that accessing ‘true’ critical behavior is not excluded per se from the neural functional. The strategy used for the extrapolation to large box sizes via the input node $1/L$ might well prove to be helpful for this purpose. Supplying additional training data that feature the slow decay of density profiles associated with long-range correlations might also be necessary. Further, the architecture of the neural network might need to be modified to accommodate the growing correlation length which could require incorporating a larger density window as input to the neural functional.
Although we have focused in this work on the prototypical Lennard-Jones fluid, for ease of comparison with previous accurate simulation data, we see no fundamental problems in applying our methodology to more complex fluids. As the formal structure of DFT remains intact for arbitrary interaction potentials, the method could be transferred directly to the investigation of systems that feature more elaborate force fields, including three- and higher-body contributions, as used e.g. to model interactions in water [88, 89] or in colloidal gels [90]. Methodological extensions arise already for the particular case of pairwise interactions only. These include the incorporation of bulk pair correlation data during training [23, 28] and the generalization of the functional dependence to feature explicitly the pair potential $\phi(r)$ [33]. In the future, it would be interesting to investigate further test particle concepts and the topical problem of inverting structural data to infer interparticle interactions, and their uniqueness [91], via neural functional methods.
For the case of anisotropic and molecular fluids, orientational degrees of freedom must be accounted for: the one-body density profile $\rho(\mathbf{r},\bm{\omega})$ which enters the functional mapping must be resolved with respect to both position $\mathbf{r}$ and orientation $\bm{\omega}$. In a first venture, Simon et al. [30] demonstrated that machine learning a classical density functional is feasible for the anisotropic Kern-Frenkel model in planar geometry. Notably, the authors found that the external flat-wall potentials for the generation of training data had to incorporate nontrivial orientation-dependence in order to probe the functional mapping sufficiently. Another very recent work considers an application to a model of carbon dioxide, where orientations become important [34]. Methods developed in the context of molecular DFT [92, 93] could serve as a practical guide on how to deal with numerical challenges arising from the additional orientational resolution; see the recent Ref. [94].
For mixtures, the relevant functionals depend on the density profiles $\rho_{i}(\mathbf{r})$ of the individual species $i=1,\dots,s$. A machine-learning scheme aiming to represent the direct correlation functional $c_{1,i}(\mathbf{r};[\rho_{1},\dots,\rho_{s}],T)$ must therefore account for the additional species labeling. Tackling the fluid phase behavior of multi-component systems is, of course, important in physical chemistry and chemical engineering. Note, for example, that rich phase behavior emerges already for the case of a very simple binary mixture [95]. In ionic systems, charge ordering is key and appropriate number-number and charge-charge correlation functions should be distinguished. For the restricted primitive model, Bui and Cox [96] have successfully demonstrated training a neural density functional, taking into account ideas from local molecular field theory to curb the long-range electrostatic interactions.
@@ Acknowledgements.
We thank Francesco Turci and Nigel Wilding for helpful discussions. R. E. acknowledges support of the Leverhulme Trust, grant no. EM-2020-029/4. This work is supported by the DFG (Deutsche Forschungsgemeinschaft) under project nos. 551294732 and 422127126.